# SPECIAL RELATIVITY

### v9

@@ 5.2 Absolute Time Coordinatization in the Accelerated Systems
We will consider the problem of aberration of light in the accelerated systems on the basis of the theory of special relativity. Let us demonstrate that for explanation of the optical effects in the rotating frame of reference one does not need neither modify the special theory of relativity, nor apply the general theory of relativity. It is only necessary to strictly follow the special theory of relativity. In order to write electrodynamics laws in the accelerated frame we have to go just one step further, and define the metric of the non-inertial frame. One must take into account that metric tensor must be a continuous quantity. The first mathematical idea is the smooth tailoring of the metric tensor. This problem in special relativity can be adequately treated only by an approach which uses the absolute time coordinatization. We describe the mathematical expansion of special theory of relativity onto accelerations in the Chapter 9. Also this can be done directly, it is some times possible to save time by getting the answer with some physical arguments. For example, in our case of interest one can take advantage of the following considerations, which are in agreement with rigorous mathematical derivation.
Suppose that the initial inertial system $S$ is at rest with respect to the fixed stars and a system $S_{n}$ (and an observer with his measuring instruments) in the frame $S$ is accelerated from the rest with respect to the fixed stars up to velocity $v$ along the $x$-axis. In accelerated systems, only the theory maintaining an absolute simultaneity is logically consistent with the natural behavior of clocks. The method of synchronization consists in keeping, without changes, the same set of uniformly synchronized clocks used in the case when the system $S_{n}$ was at rest. It is well known that during the motion with acceleration (with respect to the fixed stars) the procedure of Einstein’s clock synchronization cannot be performed and the interval in the accelerated reference system $S_{n}$ will, by the moment when the system $S_{n}$ starts moving with constant velocity, have the non-diagonal form.
Absolute simultaneity can be introduced in special relativity without affecting neither the logical structure, no the (convention-independent) predictions of the theory. We begin with the metric as the true measure of space-time intervals for an non-inertial observer $S_{n}$ with coordinates $(t_{n},x_{n},y_{n},z_{n})$. We transform coordinates $(t,x,y,z)$ that would be coordinates of an inertial observer $S$ moving with velocity $-v$ with respect to the observer $S_{n}$, using the inverse Galilean transformation: we substitute $x=x_{n}+vt$, while leaving time unchanged $t_{n}=t$ into the Minkowski metric $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$ to obtain (1)
@@
This metric describes the electrodynamics of the accelerated light source with viewpoint of of the accelerated observer measurements as viewing this of the inertial observer. Inspecting Eq. (15) we can find the components of the metric tensor $g_{\mu\nu}$ in the coordinate system of $S_{n}$. We obtain $g_{00}=1-v^{2}/c^{2}$, $g_{01}=-v/c$, $g_{11}=-1$. Note that the metric in Eq. (15) is not diagonal, since, $g_{01}\neq 0$, and this implies that time is not orthogonal to space. This Langevin metric is found by matching the accelerated frame and (initial) inertial frame metric tensors. Taking into account the Langevin metric we obtain that the length of the physical rod $dx$ and the physical (proper) time $dt$ in the inertial frame will coincide with the coordinate distance and the coordinate time in the accelerated frame, i.e. $dx_{n}=dx$, $dt_{n}=dt$. (see Chapter 9 for more details).
The velocity of light emitted by a source at rest in the coordinate system $(t,x)$ for $S$ is $c$. In that case the Minkowski metric Eq.(1) associated with inertial frame $S$ predict a symmetry in the one-way speed of light. In the coordinate system $(t_{n},x_{n})$, however, the speed of light emitted by the accelerated source (i.e. source which is at rest with respect to the accelerated frame $S_{n}$) cannot be equal $c$ anymore because $(t_{n},x_{n})$ is related to $(t,x)$ via a Galilean transformation. This is readily verified if one recalls that the velocity of light in the reference system $S$ is equal to $c$. If $ds$ is the infinitesimal displacement along the world line of a ray of light, then $ds^{2}=0$ and we obtain $c^{2}=(dx/dt)^{2}$. In the accelerated reference system, since $x_{n}=x-vt$ and $t=t_{n}$, this expression takes the form $c^{2}=(dx_{n}/dt_{n}+v)^{2}$, which can be seen by setting $ds^{2}=0$ in Eq. (15). This means that in the accelerated reference system of coordinates $(ct_{n},x_{n})$ the velocity of light parallel to the x-axis, is $dx_{n}/dt_{n}=c-v$ in the positive direction, and $dx_{n}/dt_{n}=-c-v$ in the negative direction as stated above. The reason why it is different from the electrodynamics constant $c$ is due to the fact that the clocks are synchronized following the absolute time convention, which is fixed because $(t_{n},x_{n})$ is related to $(t,x)$ via a Galilean transformation.
@@ 5.3 The Asymmetry between the Inertial and Accelerated Frames
We discovery of the essential asymmetry between the inertial and accelerated frames, namely, the Maxwell’s equations are not applicable from the viewpoint of an observer at rest with respect to accelerated frame $S_{n}$. In fact, the metric Eq.(15) associated with accelerated reference frame $S_{n}$ predicts an asymmetry in the one-way speed of light in the relative velocity direction. Accelerations (with respect to the fixed stars) have an effect on the propagation of light. On accelerated system $S_{n}$, the velocity of light emitted by a source at rest must be added to (or subtracted from) the speed due to acceleration and the velocity of light is different in opposite directions. In contrast, the Maxwell’s equations continue to hold from the viewpoint of an observer at rest with respect to inertial frame $S$. On inertial system $S$, the velocity of light emitted by a source at rest is $c$.
@@
The following important detail of such acceleration frame description can hardly be emphasize enough. Metric applies to physical laws, not to physical facts. Indeed, it is always possible to chose such variables, in which metric of the accelerated source will be diagonal. We interpret the Langevin metric to mean that the law of electrodynamics is expressed by anisotropic field equations in the accelerated frame. By a physical facts in this context we mean the aberration of light radiated by a single ”plane wave” emitter in the accelerated frame. The electrodynamics equation needs to be integrated with initial condition for the radiation wavefront. After the boost we can see that acceleration has no effect on the wavefront orientation. In fact, the variables $(t,x,y,z)$ can be expressed in terms of the variables $(t_{n},x_{n},y_{n},z_{n})$ by means of Galilean transformation $x_{n}=x-vt$, $t_{n}=t$, so that the wavefront of the emitted light beam is perpendicular to the vertical direction $z_{n}$ after the acceleration, Fig. 14. We shall discuss this subject further in the Section 5.5.
It would be well to begin with bird’s view of some main results. Suppose that an observer in the accelerated frame $S_{n}$ performs an aberration measurement. How shall we describe the aberration of light from the ”plane wave” emitter which is at rest in the inertial frame $S$? In order to predict the result of the aberration measurement the accelerated observer should use the non-diagonal metric Eq.(15).
@@
Figure 14: Aberration of light in an accelerated frame of reference $S_{n}$. Radiation wavefront orientation and the anisotropy of speed of light presented in the absolute time coordinatization ($t_{n}=t$) for the screen. The crossed term in metric Eq.(15) generates anisotropy in the accelerated frame that is responsible for the change of radiation direction (aberration).
@@
Figure 15: Aberration of light in an accelerated frame of reference $S_{n}$. Screen is at rest in the accelerated frame. Radiation wavefront orientation and the anisotropy of speed of light presented in the absolute time coordinatization ($t_{n}=t$) for the screen. The aberration of the transmitted light is considered independent of the source speed and to have just a local origin exclusively based on the observer acceleration relative to the fixed stars.
@@
According to the asymmetry between the inertial and accelerated frames, there is a remarkable prediction on the theory of the aberration of light. Namely, if the opaque screen with hole was at rest relative to the fixed stars and the screen started from rest to motion, then the apparent angular position of the ”plane-wave” emitter seen in the accelerated frame through the aperture would jump by angle $-v/c$. That is the transmission through the hole in the opaque screen in the frame $S_{n}$ appears as shown in Fig 14. It is important to note that the source of asymmetry is not the relative acceleration of two frames; rather, it is the difference in the accelerational history of each frame separately with respect to the fixed stars.
@@
Above (see Chapter 4) we considered the emitter-screen problem in the frame of reference $S$ which is at rest (i.e. without accelerational history) with respect to the fixed stars. It is found that there is no aberration proceeding from the emitter, independently of their motion in the case of the small aperture hole. It is important to emphasize that the aberration of the light beam transmitted through the small aperture hole also does not dependent on the emitter motion in the accelerated system $S_{n}$, Fig. 15. The point is that the crossed term in metric Eq.(15), which generates aberration in this particular case, depends only on the velocity change of the accelerated frame relative to the initial inertial frame.
@@ 5.4 Explanation of the Aberration on the Basis of Electrodynamics
It is well known that the electrodynamics theory meets all requirements of the theory of relativity and therefore must accurately describe the properties of such a relativistic object as light. For example, we are able to demonstrate that velocity of light in the accelerated frame can be derived independently of a technique involving metric Eq. (15). In the inertial frame, fields are expressed as a function of the independent variables $x,y,z$, and $t$. According to the special relativity, the Maxwell’s equations always valid in the Lorentz reference frame. The electric field $\vec{E}$ of an electromagnetic wave radiated by the emitter at rest in the inertial frame satisfies the equation $\Box^{2}\vec{E}=\nabla^{2}\vec{E}-\partial^{2}\vec{E}/\partial(ct)^{2}=0$. However, the variables $x,y,z,t$ can be expressed in terms of the variables $x_{n},y_{n},z_{n},t_{n}$ by means of a Galilean transformation, so that fields can be written in terms of $x_{n},y_{n},z_{n},t_{n}$. Hence the wave equation transforms into
@@
where coordinates and time are transformed according to a Galilean transformation. The solution of this equation $F[x_{n}-(c-v)t_{n}]+G[x_{n}+(-c-v)t_{n}]$ is the sum of two arbitrary functions, one of argument $x-(c-v)t$ and the other of argument $x+(-c-v)t$. The first term represents a wave traveling forward in the positive $x_{n}$ direction, and the second term a wave traveling backwards in the negative $x_{n}$ direction. Acceleration has an effect on the propagation of light. In fact, the coordinate velocity of light parallel to the $x$-axis is given by $dx/dt=c-v$ in the positive direction, and $dx/dt=-c-v$ in the negative direction.
We now consider the transmission through the hole in the opaque screen in the accelerated frame $S_{n}$, Fig. 15. We describe the aberration of light based on the observations made by an observer in the same accelerated frame as the screen. It is assumed that the detector for the direction of the radiation is an energy propagation detector and the size of the detector aperture is sufficiently large compared with the radiation beam size. In other words, it is assumed that there is no physical influence of the detector (e.g. aperture) on the measurement.
We have already discussed the explanation of the aberration on the basis of electrodynamics in the previous chapter. Our earlier discussion is really about as far as anyone would normally need to go with the subject, but we are going to do it again. One reason is that one should know how to deal with what happens to waves in the accelerated frame from the viewpoint electrodynamics equations. First, we remind about the convenient way of describing diffraction by opaque screen we used in the previous chapter. The screen containing a hole is a kind of diffraction grating which breaks up the radiated beam into a number of diffracted beams of plane waves. Each of these beams corresponds to one of the Fourier components into which an transmittance can be resolved. Let us assume that the transmittance of the grating varies according to the law $T=g(K_{\perp})\cos{({K}_{\perp}x)}$.
@@
In our example, the plane waves with different wavenumber vectors propagate out from the screen with different frequencies. Then equation $\Delta\omega/\Delta k_{x}=-v$ holds for each transmitted waves independently on the sign and the magnitude of the transmitted angle. The last equations state that transmitted light beam with finite transverse size moves along the $x$ direction with group velocity $d\omega/dk_{x}=-v$, Fig. 15.
@@
We would now like to discuss an apparent paradox. We obtained the electrodynamics equations in accelerated frame using the Galilean transformation of the Maxwell’s equations. Intuition would seem to tell us that everything is at rest in the accelerated frame. But special relativity says that there is a derivative $\partial/\partial t_{n}=v\partial/\partial x_{n}$ because there is $\partial/\partial x_{n}$ that is not zero. Here we have a new kind of situation which is quite different from inertial frame electrodynamics. We will discuss this subject further in the Chapter 10.

### v10

@@ 5.1 Absolute Time Coordinatization in the Accelerated Systems
We will consider the problem of aberration of light in accelerated systems on the basis of the theory of special relativity. Let us demonstrate that for an explanation of the optical effects in the rotating frame of reference, one does not need to modify the special theory of relativity nor apply the general theory of relativity. It is only necessary to strictly follow the special theory of relativity. To write electrodynamics laws in the accelerated frame we have to go just one step further and define the metric of the non-inertial frame. One must take into account that the metric tensor must be a continuous quantity. The first mathematical idea is the smooth tailoring of the metric tensor. This problem in special relativity can be adequately treated by an approach that uses absolute time coordinatization. We describe the mathematical expansion of the special theory of relativity onto accelerations at the end of this subsection. Also, this can be done directly, it is sometimes possible to save time by getting the answer with some physical arguments. For example, in our case of interest, one can take advantage of the following considerations, which are in agreement with rigorous mathematical derivation.
Suppose that the initial inertial system $S$ is at rest with respect to the fixed stars and a system $S_{n}$ (and an observer with his measuring instruments) in the frame $S$ is accelerated from the rest with respect to the fixed stars up to velocity $v$ along the $x$-axis. In accelerated systems, the theory of maintaining absolute simultaneity is logically consistent with the smooth tailoring of the metric tensor. The method of synchronization consists in keeping, without changes, the same set of uniformly synchronized clocks used in the case when the system $S_{n}$ was at rest. It is well known that during the motion with acceleration (with respect to the fixed stars) the procedure of Einstein’s clock synchronization cannot be performed and the interval in the accelerated reference system $S_{n}$ will, by the moment when the system $S_{n}$ starts moving with constant velocity, have the non-diagonal form.
We begin with the metric as the true measure of space-time intervals for an accelerated observer $S_{n}$ with coordinates $(t_{n},x_{n},y_{n},z_{n})$. We transform coordinates $(t,x,y,z)$ that would be coordinates of an inertial observer $S$ moving with velocity $-v$ with respect to the observer $S_{n}$, using the inverse Galilean transformation: we substitute $x=x_{n}+vt$, while leaving time unchanged $t_{n}=t$ into the Minkowski metric $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$ to obtain
@@
This metric describes the electrodynamics of the accelerated light source with the viewpoint of the accelerated observer measurements as viewing this of the inertial observer. Inspecting Eq.(15) we can find the components of the metric tensor $g_{\mu\nu}$ in the coordinate system of $S_{n}$. We obtain $g_{00}=1-v^{2}/c^{2}$, $g_{01}=-v/c$, $g_{11}=-1$. Note that the metric in Eq. (15) is not diagonal, since, $g_{01}\neq 0$, and this implies that time is not orthogonal to space. This Langevin metric is found by matching the accelerated frame and (initial) inertial frame metric tensors. Taking into account the Langevin metric we obtain that the length of the physical rod $dx$ and the physical (proper) time $dt$ in the inertial frame will coincide with the coordinate distance and the coordinate time in the accelerated frame, i.e. $dx_{n}=dx$, $dt_{n}=dt$. (see Chapter 12 for more details).
Now let us return to the mathematical expansion of the special theory of relativity onto accelerations. We will keep the mathematical complexity to a minimum and assume that the frame $S_{n}$ begins at $t=0$ to move with constant acceleration $a>0$ and does so till $t=T$. Now we would like to discuss an expression for the velocity from the viewpoint of the inertial (non-accelerated) observer. We know that uniformly accelerated motion must obey the equation $d(\gamma\vec{v})/dt=\vec{f}/m=\vec{a}$. This is readily integrated $\gamma\vec{v}=\vec{a}t$. We will then have $dx/dt=at/\sqrt{1+a^{2}t^{2}/c^{2}}$. We emphasize that $dx(t)/dt$ is the velocity of the noninertial frame as seen by the inertial observer. Under the inverse Galilean transformation $dx=dx_{n}+atdt/\sqrt{1+a^{2}t^{2}/c^{2}}$, $dt_{n}=dt$ the metric of the uniformly accelerating observer becomes
Since the metric tensor of space-time must be a continuous quantity, we must match the coordinates in the accelerated and inertial stages. At $t=0$ this metric naturally coincides with the $S$ frame metric and we have $g_{00}=1$. During acceleration, we picked the absolute time coordinatization. Therefore, for the metric tensor to be continuous when the reference frame passes from non-inertial motion to inertial one we will make use of the metric Eq. 16. We will therefore require that at $t=T$ the metric of a noninertial frame would pass continuously into the Langevin metric. For this purpose, it is sufficient to relate the coordinates and time of the accelerated observer $(x_{n},t_{n})$ to the coordinates $(x,t)$ of the inertial observer by the Galilean boost $x_{n}=x-vt$, $t_{n}=t$, where $v=aT/\sqrt{1+a^{2}T^{2}/c^{2}}$. In this case, in the interval $T<t_{n}$ of coordinate time $t_{n}$ the reference frame of the accelerated system has the metric Eq. (15). We emphasize that $v$ is the velocity of the frame $S_{n}$ as seen by the inertial observer at the second (inertial) stage of $S_{n}$ motion. It becomes obvious that at $t=t_{n}=T$ the metric of the noninertial frame goes into the metric of the (accelerated) inertial frame in a continuous manner [^25] For a general discussion of the mathematical expansion of special relativity onto acceleration we suggest reading the book [16]..
According to textbooks, inertial frames are characterized by motion at constant velocity relative to the fixed stars, and different inertial frames are related by Lorentz boosts. Therefore, it might appear that the coordinates and time in frames $S$ and $S_{n}$ (in the inertial stage) should be related by the Lorentz transformation. But here at the inertial part of the trajectory, the metric of both frames will be diagonal in the Lorentz coordinatization and so at $t_{n}=T$, the metric tensor of the accelerated frame must abruptly change from the value $g_{00}=(1+a^{2}T^{2}/c^{2})^{-1}$ to the value $g_{00}=1$.
@@ 5.2 The Asymmetry between the Inertial and Accelerated Frames
The velocity of light emitted by a source at rest in the coordinate system $(t,x,y,z)$ for $S$ is $c$. In that case the Minkowski metric Eq.(1) associated with inertial frame $S$ predicts a symmetry in the one-way speed of light. In the coordinate system $(t_{n},x_{n},y_{n},z_{n})$, however, the speed of light emitted by the accelerated source (i.e. source which is at rest with respect to the accelerated frame $S_{n}$) cannot be equal $c$ anymore because $(t_{n},x_{n},y_{n},z_{n})$ is related to $(t,x,y,z)$ via a Galilean transformation. This is readily verified if one recalls that the velocity of light in the reference system $S$ is equal to $c$. If $ds$ is the infinitesimal displacement along the world line of a ray of light, then $ds^{2}=0$ and we obtain $c^{2}=(dx/dt)^{2}$. In the accelerated reference system, since $x_{n}=x-vt$ and $t=t_{n}$, this expression takes the form $c^{2}=(dx_{n}/dt_{n}+v)^{2}$, which can be seen by setting $ds^{2}=0$ in Eq. (15). This means that in the accelerated reference system of coordinates $(ct_{n},x_{n})$ the velocity of light parallel to the x-axis, is $dx_{n}/dt_{n}=c-v$ in the positive direction, and $dx_{n}/dt_{n}=-c-v$ in the negative direction as stated above.
@@
We discovered the essential asymmetry between the inertial and accelerated frames, namely, Maxwell’s equations are not applicable from the viewpoint of an observer at rest with respect to accelerated frame $S_{n}$. The metric Eq.(15) associated with accelerated reference frame $S_{n}$ predicts an asymmetry in the one-way speed of light in the relative velocity direction. Accelerations (with respect to the fixed stars) affect on the propagation of light. On accelerated system $S_{n}$, the velocity of light emitted by a source at rest must be added to (or subtracted from) the speed due to acceleration, and the velocity of light is different in opposite directions. In contrast, on an inertial system, $S$, the velocity of light emitted by a source at rest is $c$.
@@
The following important detail of such acceleration frame description can hardly be emphasized enough. Metric applies to physical laws, not to physical facts. Indeed, it is always possible to choose such variables, in which the metric of the accelerated source will be diagonal. We interpret the Langevin metric to mean that the law of electrodynamics is expressed by anisotropic field equations in the accelerated frame. By physical facts in this context we mean the aberration of light radiated by a single ”plane wave” emitter in the accelerated frame. The electrodynamics equation needs to be integrated with the initial condition for the radiation wavefront. After the boost we can see that acceleration does not affect the wavefront orientation. In fact, the variables $(t,x,y,z)$ can be expressed in terms of the variables $(t_{n},x_{n},y_{n},z_{n})$ by means of Galilean transformation $x_{n}=x-vt$, $t_{n}=t$, so that the wavefront of the emitted light beam is perpendicular to the vertical direction $z_{n}$ after the acceleration, Fig. 14. We shall discuss this subject further in Section 5.5.
@@
Figure 14: Aberration of light in an accelerated frame of reference $S_{n}$. Radiation wavefront orientation and the anisotropy of the speed of light are presented in the absolute time coordinatization ($t_{n}=t$) for the screen.
@@
Figure 15: Aberration of light in an accelerated frame of reference $S_{n}$. The screen is at rest in the accelerated frame. The aberration of the transmitted light is considered independent of the source speed.
@@
It would be well to begin with a bird’s view of some main results. Suppose that an observer in the accelerated frame $S_{n}$ performs an aberration measurement. To predict the result of the aberration measurement the accelerated observer should use the non-diagonal metric Eq.(15). According to the asymmetry between the inertial and accelerated frames, there is a remarkable prediction on the theory of the aberration of light. Namely, if the opaque screen with a hole was at rest relative to the fixed stars and the screen started from rest to motion, then the apparent angular position of the ”plane-wave” emitter seen in the accelerated frame through the aperture would jump by angle $-v/c$. That is the transmission through the hole in the opaque screen in the frame $S_{n}$ appears as shown in Fig 14. It is important to note that the source of asymmetry is not the relative acceleration of two frames; rather, it is the difference in the accelerational history of each frame separately with respect to the fixed stars.
@@
Above (see Chapter 4) we considered the emitter-screen problem in the frame of reference $S$ which is at rest (i.e. without accelerational history) with respect to the fixed stars. It is found that there is no aberration proceeding from the emitter, independently of their motion in the case of the small aperture hole. It is important to emphasize that the aberration of the light beam transmitted through the small aperture hole also is not dependent on the emitter motion in the accelerated system $S_{n}$, Fig. 15. The point is that the crossed term in metric Eq.(15), which generates aberration in this particular case, depends only on the velocity change of the accelerated frame relative to the initial inertial frame.
@@ 5.3 Explanation of the Aberration on the Basis of Electrodynamics
It is well known that the electrodynamics theory meets all requirements of the theory of relativity. For example, we can demonstrate that the velocity of light in the accelerated frame can be derived independently of a technique involving metric Eq. (15). In the inertial frame, fields are expressed as a function of the independent variables $x,y,z$, and $t$. According to the special relativity, Maxwell’s equations are always valid in the Lorentz reference frame. The electric field $\vec{E}$ of an electromagnetic wave radiated by the emitter at rest in the inertial frame satisfies the equation $\Box^{2}\vec{E}=\nabla^{2}\vec{E}-\partial^{2}\vec{E}/\partial(ct)^{2}=0$. However, the variables $x,y,z,t$ can be expressed in terms of the variables $x_{n},y_{n},z_{n},t_{n}$ using a Galilean transformation, so that fields of an electromagnetic wave radiated by the emitter at rest in the accelerated frame can be written in terms of $x_{n},y_{n},z_{n},t_{n}$. Hence the wave equation transforms into
@@
where coordinates and time are transformed according to a Galilean transformation. The solution of this equation $F[x_{n}-(c-v)t_{n}]+G[x_{n}+(-c-v)t_{n}]$ is the sum of two arbitrary functions, one of argument $x-(c-v)t$ and the other of argument $x+(-c-v)t$. The first term represents a wave traveling forward in the positive $x_{n}$ direction, and the second term a wave traveling backward in the negative $x_{n}$ direction. Acceleration affects the propagation of light. The coordinate velocity of light parallel to the $x$-axis is given by $dx/dt=c-v$ in the positive direction, and $dx/dt=-c-v$ in the negative direction.
We now consider the transmission through the hole in the opaque screen in the accelerated frame $S_{n}$, Fig. 15. We describe the aberration of light based on the observations made by an observer in the same accelerated frame as the screen. We have already discussed the explanation of the aberration based on electrodynamics. Our earlier discussion is really about as far as anyone would normally need to go with the subject, but we are going to do it again. One reason is that one should know how to deal with what happens to waves in the accelerated frame from the viewpoint of electrodynamics equations. The screen containing a hole is a kind of diffraction grating that breaks up the radiated beam into a number of diffracted beams of plane waves. Each of these beams corresponds to one of the Fourier components into which a transmittance can be resolved. Let us assume that the transmittance of the grating varies according to the law $T=g(K_{\perp})\cos{({K}_{\perp}x)}$.
@@
In our example, the plane waves with different wavenumber vectors propagate out from the screen with different frequencies. Then equation $\Delta\omega/\Delta k_{x}=-v$ holds for each transmitted wave independently on the sign and the magnitude of the transmitted angle. The last equations state that the transmitted light beam with finite transverse size moves along the $x$ direction with group velocity $d\omega/dk_{x}=-v$, Fig. 15.
@@
We would now like to discuss an apparent paradox. We obtained the electrodynamics equations in an accelerated frame using the Galilean transformation of Maxwell’s equations. Intuition would seem to tell us that everything is at rest in the accelerated frame. But special relativity says that there is a derivative $\partial/\partial t_{n}=v\partial/\partial x_{n}$ because there is $\partial/\partial x_{n}$ that is not zero. Here we have a new kind of situation that is quite different from inertial frame electrodynamics. We will discuss this subject further in Chapter 9.
5.4 Simple Explanation of the Aberration of Light Effect in Accelerated Frame
Above we considered the aberration of light problem in the accelerated frame. The aberration of light effect is not easy to calculate in the frame of reference attending acceleration. In this case, we used a metric tensor to obtain the electrodynamics equations in the accelerated reference frame. The presented approach uses the Fourier transform method. As one of the consequences of the crossed term in the wave equation, we found the group velocity that is responsible for the aberration. Now let us present a very simple explanation of the aberration of light in the accelerated frame based on electrodynamics. To find the aberration increment, we simply use Galilean transformation law for the electric and magnetic fields.
We have an expression for the energy flow vector of the electromagnetic field. This vector, $\vec{S}=c\vec{E}\times\vec{B}/(4\pi)$, is called ”Poynting’s vector”. It tells us the rate at which the field energy moves around in space. To detect the aberration of light effect inside the accelerated frame, it is obvious that some coordinate system with reference direction is needed. We determine the reference directions perpendicular to the line motion in the inertial frame and in the accelerated frame. The motion of the aberrated light beam is assumed for simplicity, to lie in the same plane and the angular position of the aberrated beam is described by one angle. The analysis is much simplified if we treat separately the case of a radiated beam with its $\vec{E}$-vector parallel to the ”plane of incidence” (that is $xz$-plane) and the case of a radiated beam with the $\vec{E}$-vector perpendicular to the $xz$-plane. We will carry through the analysis for an incoming beam polarized perpendicular to the plane of incidence, but the principle is the same for both. So we take that $\vec{E}$, has only $y$- component. For a magnetic field, we get $\vec{B}=\vec{e}_{z}\times\vec{E}$, where $\vec{e}_{z}$ is the direction of Poynting vector $\vec{S}$ in the inertial frame.
Above we demonstrated that we must match the (4D) coordinates in the accelerational and inertial parts of the trajectory. For this purpose, it is sufficient to relate the coordinates and time of the accelerated observer to the coordinates and time of the inertial observer by the Galilean boost. We can obtain the electrodynamics equations in an accelerated frame using the Galilean transformation (with velocity $-v$) of Maxwell’s equations. The fields $\vec{E}_{n}$ and $\vec{B}_{n}$ observed in the accelerated frame where the emitter is at rest are not the same as those $\vec{E}$ and $\vec{B}$ in the inertial frame before the active Galilean boost. With the Galilean transformation, the field transformation is $\vec{E}_{n}=\vec{E}-\vec{v}\times\vec{B}/c$, $\vec{B}_{n}=\vec{B}-\vec{v}\times\vec{E}/c$. This expression describes the first-order (in $v/c$) effects only. Here $\vec{v}=v\vec{e}_{x}$ is the velocity vector of the accelerated frame in the inertial frame. If the $\vec{E}$-vector perpendicular to the $xz$-plane, the magnetic field in the inertial frame has only a $x$-component and we have immediately $\vec{E}_{n}=\vec{E}=\vec{e}_{y}E$, $\vec{B}_{n}=(\vec{e}_{z}\times\vec{e}_{y})E-(\vec{e}_{x}\times\vec{e}_{y})vE/c=-\vec{e}_{x}E-\vec{e}_{z}vE/c$. Therefore there is a Poynting vector, $c\vec{E}_{n}\times\vec{B}_{n}/(4\pi)=[-(\vec{e}_{y}\times\vec{e}_{x})-(\vec{e}_{y}\times\vec{e}_{z})v/c]cE^{2}/(4\pi)=[\vec{e}_{z}-\vec{e}_{x}(v/c)]cE^{2}/(4\pi)$. Then, the radiated beam is propagated at the angle $-v/c$ with respect to the $z_{n}$-axis yielding the phenomenon of the aberration of light in the accelerated frame.
We have found that we get the same result whether we analyze the energy transport direction using electrodynamics equations in the accelerated frame or using Galilean transformation law for electromagnetic fields. In the first instance, the equations state that the transmitted light beam with finite transverse size moves along $x$-direction with group velocity $d\omega/dk_{x}=-v$, in the second, the effect of light aberration is understood as a change in the direction of Poynting vector. The two approaches give the same result. In this section, we demonstrate that the direction of group velocity and the direction of the Poynting vector are two sides of the same coin.
Acceleration affects the field equations. When the observer is at rest in the accelerated frame the derivative $\partial/\partial t_{n}$ will be modified. One can easily show this by transforming coordinates $(t,x,y,z)$ that would be coordinates of an inertial observer moving with velocity $-v$ with respect to the accelerated observer. From the Galilean transformation $x_{n}=x-vt,~{}y_{n}=y,~{}z_{n}=z,~{}t_{n}=t$, after partial differentiation, one obtains $\partial/{\partial t_{n}}=\partial/{\partial t}+v\partial/{\partial x}$, $\partial/{\partial x_{n}}=\partial/{\partial x}$. Our problem now is to work out the magnetic field of the transmitted light beam in terms of those of the incident plane wave. How can we do that? The magnetic field $\vec{B}_{n}$ satisfies electrodynamics equation ×$\vec{\nabla}\times\vec{B}_{n}=(\partial/{\partial t}+v\partial/{\partial x})\vec{E}_{n}/c$, in the accelerated frame. We are very specific and wright out explicitly the new component: $-\partial B_{z}/{\partial x}=(v/c)\partial E_{y}/{\partial x}$. In any measurement of the angular displacement of transmitted (through aperture) light beam, one will have to consider a finite radiation beam size. If we integrate this equation with respect to $x$ across transmitted light beam, we conclude that $\vec{B}_{n}=-\vec{e}_{x}E-\vec{e}_{z}vE/c$. This last result is just what we got by a field transformation argument.
This is a good point to make a general remark about the physical reality of the energy flow (Poynting) vector. In any physically realizable aberration measurement, one will have to consider a finite radiation beam size. The direction of the energy transport (and, hence, the direction of the Poynting vector) has an exact objective meaning i.e. convention-invariant. However, there is a common mistake made in electrodynamics connected with the energy transport direction in the case of a plane wave. It is impossible to know the energy transport direction when one has to deal with a plane wave.